OpenStudy (anonymous):

Which of the following quadrilaterals is not a parallelogram? A. square B. rectangle C. rhombus D. trapezoid

OpenStudy (goldphenoix):

Trapezoid.

OpenStudy (goldphenoix):

The definition of a parallelogram is " a parallelogram is a simple (non self-intersecting) quadrilateral with two pairs of parallel sides. The opposite or facing sides of a parallelogram are of equal length and the opposite angles of a parallelogram are of equal measure. "

OpenStudy (anonymous):

that's what I thought thanks for your help

OpenStudy (goldphenoix):

No problem. I hope this help!

OpenStudy (dan815):

parallelogram is where u got a gram of parallels

OpenStudy (dan815):

diagram of parallel lines

OpenStudy (anonymous):

can anyone help me with this problem? A parallelogram has an area of 96 cm2 and a height of 4 cm. What is the base length of the parallelogram? Explain your steps to solving the problem.

jimthompson5910 (jim_thompson5910):

Area = base x height A = b*h 96 = b*4 solve for b

OpenStudy (dan815):